### v1

@@ Abstract
This paper presents a novel identification approach of Koopman models of nonlinear systems with inputs under rather general noise conditions. The method uses deep state-space encoders based on the concept of state reconstructability and an efficient multiple-shooting formulation of the squared loss of the prediction error to estimate the dynamics and the lifted state from input-output data. Furthermore, the Koopman model structure includes an innovation noise term that is used to handle process and measurement noise. It is shown that the proposed approach is statistically consistent and computationally efficient due to the multiple-shooting formulation where, on subsections of the data, multi-step prediction errors can be calculated in parallel. The latter allows for efficient batch optimization of the network parameters and, at the same time, excellent long-term prediction capabilities of the obtained models. The performance of the approach is illustrated by nonlinear benchmark examples.
@@ 1 Introduction
Due to the continuously increasing performance expectations for dynamical systems in engineering, nonlinear behavior in many application areas started to become dominant, requiring novel methods that can stabilize and shape the performance of these systems with the ease of conventional approaches that have been developed for linear time-invariant systems. Hence, recent years have seen a strong push to find global linear embeddings of nonlinear systems to simplify, among others, analysis, prediction and control. One such embedding technique is based on the Koopman framework, where the concept is to lift the nonlinear state space to a (possibly) infinite-dimensional space through the so-called observable functions. The dynamics of the original system are preserved and governed by a linear Koopman operator, enabling the representation of the system dynamics via a linear dynamical description [6], [28]. In practice, by choosing a dictionary of a finite number of observables a priori to construct time-shifted data matrices, linear Koopman-based models have been commonly obtained using simple least squares estimation [27]. One such approach, called dynamic mode decomposition (DMD) [38], is based on constructing time-shifted data matrices using directly measured state variables associated with the system. If the dictionary consists of nonlinear functions of the state, this technique is known as extended DMD (EDMD) [52]. However, among many challenges related to statistical consistency, availability of state-measurements, etc., the main difficulty with these powerful methods lies in choosing a finite number of lifting functions such that, in the lifted state-space, a linear time-invariant (LTI) model exists that can capture well the dynamic behavior of the original nonlinear system. While there exist methods for the automatic selection of the observables (see [7], [54]), they still rely on an a priori choice of a dictionary of functions, which many times are difficult to select and even characterize the resulting approximation error by them.
@@
To circumvent this, a viable approach has been found in learning the lifting functions from data by the use of machine learning methods such as Guassian processes [20], kernel-based methods [36], [53], or various forms of Artificial Neural Networks (ANNs) [24, 34, 35, 50]. Due to their flexibility in describing multiple model structures, applicability to large datasets, many successful applications of these methods have been reported in the literature to obtain accurate and compact Koopman models in practice.
@@
However, a common drawback of learning-based methods together with the (E)DMD approaches is (i) the assumption of full-state measurement (e.g. [24], [35]), which is rarely the case in engineering applications. Some works such as [17] and [54] do address partial state observations, either by only lifting the output [54] or by implementing a DMD version that uses time-delayed measurements [17]. In a different approach, [34] employs a Kalman filter to estimate the lifted state. Nevertheless, a systematic framework for addressing partial state measurements in the context of data-driven Koopman modelling is still lacking at large. Furthermore, despite the powerful capabilities of these approaches that have been demonstrated in multiple examples, generally (ii) little consideration of measurement or process noise is taken, which can lead to serious bias of the models when applied in real-world applications. Only a few papers present examples where measurement noise is even present in the data (e.g. [15], [48]) and often only the robustness of the methods is analyzed (e.g. [43]). While there are works that add process noise directly to the lifted representation (e.g. [34]), the way noise enters the Koopman model is merely an assumption. As such, ensuring statistical guarantees of consistency of the estimators remains an open question in the literature. A third important issue is that (iii) the estimation of Koopman models for systems with inputs has only recently been investigated, either through a nonlinear lifting [3] or by using state- and input-dependent observables, together with input increments [50]. However, this often leads to models that are not directly useful for engineering applications, as without direct use of inputs, it is difficult to analyze dynamical aspects of the system or to design controllers to regulate the behavior compared to other model classes. Due to their simple structure, multiple works assume a fully LTI Koopman model (e.g. [19], [26]) or, lately, bilinear (e.g. [4, 34, 41]), however how the approximation capability of these model structures in a finite dimensional setting compares to using more complex input functions, as given in [9, 12, 42], is still largely unexplored.
@@
To overcome challenges (i)-(iii), we introduce a flexible Koopman model learning method under control inputs, partial measurements, and with statistical guarantees of consistency under process and measurement noise. For this purpose, a deep-learning-based state-space encoder approach is proposed, which is implemented in the deepSI toolbox111deepSI toolbox available at https://github.com/MaartenSchoukens/deepSI in Python. The main advantages of the approach together with our contributions [^2] The present paper extends the conference paper [9] in terms of introducing an innovation noise structure in the Koopman model to handle process and measurement noise, proving the consistency of the estimator, studying various lifted structures for control inputs and providing extensive analysis and testing of the capabilities of the method on benchmarks and real-world data. are as follows:
@@ 2.1 Koopman embedding of nonlinear systems
with $x_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ being the state variable, $f:\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{x}}}$ is a bounded nonlinear state transition map and $k\in\mathbb{Z}$ is the discrete time. The initial condition is denoted by $x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$ and we assume that $\mathbb{X}$ is forward invariant under $f(.)$, i.e., $f(\mathbb{X})\subseteq\mathbb{X}$, see [12]. The Koopman framework uses observable functions $\phi\in\mathcal{F}$ to lift the system (1) to a higher dimensional space with linear dynamics. These observables $\phi:\mathbb{X}\rightarrow\mathbb{R}$ are scalar functions (generally nonlinear) and are from a Banach function space $\mathcal{F}$. As described in [28], the Koopman operator $\mathcal{K}:\mathcal{F}\rightarrow\mathcal{F}$ associated with (1) is defined through:
@@
Although the Koopman framework typically requires $\mathcal{F}$ to be spanned by an infinite number of basis functions to fully describe the dynamics of (1), for practical use, an $n_{\mathrm{f}}$-dimensional linear subspace $\mathcal{F}_{n_{\mathrm{f}}}\subset\mathcal{F}$ is considered, with $\mathcal{F}_{n_{\mathrm{f}}}=\mathrm{span}\left\{\phi_{j}\right\}^{n_{\mathrm{f}}}_{j=1}$. The finite-dimensional approximation of the Koopman operator $\mathcal{K}$ can be described using the projection operator $\Pi:\mathcal{F}\rightarrow\mathcal{F}_{n_{\mathrm{f}}}$, and is given by:
(4) $\mathcal{K}_{n_{\mathrm{f}}}=\left.\Pi\mathcal{K}\right|_{\mathcal{F}_{n_{\mathrm{f}}}}:\mathcal{F}_{n_{\mathrm{f}}}\rightarrow\mathcal{F}_{n_{\mathrm{f}}}.$
In practice, the Koopman matrix representation $A\in\mathbb{R}^{n_{\mathrm{f}}\times n_{\mathrm{f}}}$ is commonly used [28]:
@@
Note that there exist classes of systems, e.g., [10], [11], that admit an exact finite dimensional embedding and $\mathcal{K}_{n_{\mathrm{f}}}$ captures exactly the effect of $\mathcal{K}$. Next, we introduce the lifted state $z_{k}=\Phi(x_{k})$, where $\Phi(x_{k})=\begin{bmatrix}\phi_{1}(x_{k})&\cdots&\phi_{n_{\mathrm{f}}}(x_{k})\end{bmatrix}^{\top}$. The lifted finite dimensional linear representation of (1) is then given by:
@@
The main challenge of the Koopman framework is the selection of the observables, including their number, to obtain a suitable approximation in terms of an appropriate norm (or an exact embedding) of the nonlinear system [28]. Additionally, it is often not clearly stated in the literature that a linear system whose dynamics are governed by the Koopman matrix $A$ is only equivalent in terms of behavior (collections of all solution trajectories) to the original nonlinear system (1) if explicit nonlinear constraints are imposed on the initial condition of the lifted state, i.e., equivalent trajectories are only part of a manifold in the extended solution space. We explore this further through a simple example.
@@ 2.2 Linear representations subject to nonlinear constraints
To illustrate the concept, we consider the following polynomial system represented by (1), similar to the one described in [5]:
@@
where $a,b,c\in\mathbb{R}$ are constant parameters and $x_{k,i}$ denotes the $i^{\text{th}}$ element of $x_{k}$. By considering solutions of (7) only on $[0,\infty)$ with initial condition $x_{0}\in\mathbb{R}^{2}$, the feasible trajectories are given by:
@@
(10) $\mathcal{B}_{\mathcal{K}}=\left\{z:\mathbb{Z}^{+}_{0}\rightarrow\mathbb{R}^{3}\mid\text{s.t. }z_{k+1}=Az_{k}\right\}.$
Note that (10) represents an unrestricted LTI behavior. It is easy to show that $\Phi(\mathcal{B})\subseteq\mathcal{B}_{\mathcal{K}}$, as any $z_{k}\in\mathcal{B}_{\mathcal{K}}$ with $z_{0}\in\mathbb{R}^{3}$ for which $z_{0,3}\neq z^{2}_{0,1}$ will not correspond to a solution of (7), i.e., $\Phi^{-1}(z_{k})=x_{k}\notin\mathcal{B}$. By introducing the constraint $\Psi:\mathbb{R}^{3}\rightarrow\mathbb{R},\Psi(z_{k})=z^{2}_{k,1}-z_{k,3}$, the solution set (10) with constraint $\Psi$ is:
(11) ^$\hat{\mathcal{B}}_{\mathcal{K}}=\left\{z:\mathbb{Z}^{+}_{0}\rightarrow\mathbb{R}^{3}\mid\text{s.t. }z_{k+1}=Az_{k},\;\Psi(z_{0})=0\right\}.$
Then, it is possible to show that $\Phi(\mathcal{B})=\hat{\mathcal{B}}_{\mathcal{K}}$ holds. To illustrate this, Fig. 1 shows the simulated trajectories of system (10) with $a=0.99$, $b=0.9$ and $c=0.9$ and the constraint $\Psi$, which we call the compliant surface. As can be seen in Fig. 1, only solutions (in green) starting on the compliant surface remain on the compliant surface and correspond to solutions (in black) of the original nonlinear system (7). Our example highlights the need for additional constraints on the Koopman form, or, as we now call it, the embedding of (1), to guarantee a bijective relationship between the solution sets.
Note that, when $x_{0}$ is known and the observable set $\Phi$ is given, this nonlinear condition on the lifted states is alternatively defined by $z_{0}=\Phi(x_{0})$. However, in an identification setting where information on $x_{0}$ is not available or only partially available, to construct a lifted model with the constrained solution set, one needs to include the $\Psi(z_{0})=0$ condition. Next, we explore observability and reconstructability of $z$ in view of our findings.
@@ 2.3 Observability and reconstructability
where $h:\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{y}}}$ is a nonlinear output map. Given $x_{0}\in\mathbb{R}^{n_{\mathrm{x}}}$, the observability map for the state $x$ associated with the nonlinear dynamics represented by (1) and (12) is:
@@
where $h\circ_{2}f(x_{0})=h\circ f\circ f(x_{0})$ and $h\circ_{n}f(x_{0})=h\circ f\circ_{n-1}f(x_{0})$ for $n>2$. Let $y^{n-1}_{0}=\begin{bmatrix}y^{\top}_{0}&\cdots&y^{\top}_{n-1}\end{bmatrix}^{\top}$. As described in [31], the representation satisfies the so-called observability rank condition at $x_{0}$ if, for $n=n_{\mathrm{x}}$, the rank of the Jacobian of $\mathcal{O}_{\mathrm{x},n}$ at $x_{0}$ is $n_{\mathrm{x}}$. If this condition is met, the representation is strongly locally observable at $\mathbb{X}_{0}$, where $\mathbb{X}_{0}$ is a neighborhood of $x_{0}$, and $\mathcal{O}_{\mathrm{x},n}$ is a diffeomorphism, i.e., it is invertible, on $\mathbb{X}_{0}$ [14]. We denote its inverse as $\Lambda_{\mathrm{x},n}:\mathbb{R}^{nn_{\mathrm{y}}}\rightarrow\mathbb{X}_{0}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$, such that
@@
For simplicity, throughout the paper, we call $\Lambda_{\mathrm{x},n}$ the observability map, used to compute the initial condition $x_{0}$ from future values of the output $y$. Conversely, the reconstructability concept is used to calculate the initial condition from past values of the output. For $n\geq 1$, it holds that
@@
(16) $x_{0}=\circ_{n-1}f(\Lambda_{\mathrm{x},n}(y^{0}_{-n+1})):=\Pi_{\mathrm{x},n}(y^{0}_{-n+1})$
for all $x_{0}\in\mathbb{X}_{0}$. The function $\Pi_{\mathrm{x},n}:\mathbb{R}^{nn_{\mathrm{y}}}\rightarrow\mathbb{X}_{0}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$ is called the reconstructability map.
@@
(17) $\begin{bmatrix}y_{0}\\ y_{1}\\ \vdots\\ y_{n-1}\\ 0\end{bmatrix}=\begin{bmatrix}\begin{pmatrix}C\\ CA\\ \vdots\\ CA^{n-1}\end{pmatrix}z_{0}\\ \Psi(z_{0})\end{bmatrix}:=\begin{bmatrix}\mathcal{O}^{\text{lin}}_{\mathrm{z},n}z_{0}\\ \Psi(z_{0})\end{bmatrix}:=\mathcal{O}_{\mathrm{z},n}(z_{0}).$
In an LTI sense, the lifted system representation would be observable if there is an $n\geq 1$ such that $\text{rank}(\mathcal{O}^{\text{lin}}_{\mathrm{z},n})=n_{\mathrm{z}}$. However, as observed in Section 2.2, it is also necessary to consider the nonlinear constraints $\Psi:\mathbb{R}^{n_{\mathrm{z}}}\rightarrow\mathbb{R}^{n_{\mathrm{c}}}$ to ensure compliance of the initial condition $z_{0}$. Hence, if there is an $n\geq 1$ such that the Jacobian of $\mathcal{O}_{\mathrm{z},n}(z_{0})$ has full rank $n_{\mathrm{z}}$, which implies that the mapping $\mathcal{O}_{\mathrm{z},n}$ is locally invertible on a neighborhood $\mathbb{Z}_{0}$ of $z_{0}$, then $z_{0}$ is uniquely determined from $y^{n-1}_{0}$ and the constraint $\Psi(z_{0})$. Then, there exists a $\Lambda_{\mathrm{z},n}:\mathbb{R}^{nn_{\mathrm{y}}}\rightarrow\mathbb{Z}_{0}\subseteq\mathbb{R}^{n_{\mathrm{z}}}$, such that

### v3

@@ Abstract
This paper presents a novel identification approach of Koopman models of nonlinear systems with inputs under rather general noise conditions. The method uses deep state-space encoders based on the concept of state reconstructability and an efficient multiple-shooting formulation of the squared loss of the prediction error to estimate the dynamics and the lifted state from input-output data. Furthermore, the Koopman model structure includes an innovation noise term that is used to handle process and measurement noise. It is shown that the proposed approach is statistically consistent and computationally efficient due to the multiple-shooting formulation where, on subsections of the data, multi-step prediction errors can be calculated in parallel. The latter allows for efficient batch optimization of the network parameters and, at the same time, excellent long-term prediction capabilities of the obtained models. The performance of the approach is illustrated by nonlinear benchmark examples and an experimental quadcopter setup.
@@ 1 Introduction
Due to the continuously increasing performance expectations for dynamical systems in engineering, nonlinear behavior in many application areas started to become dominant, requiring novel methods that can stabilize and shape the performance of these systems with the ease of conventional approaches that have been developed for linear time-invariant systems. Hence, recent years have seen a strong push to find global linear embeddings of nonlinear systems to simplify, among others, analysis, prediction and control. One such embedding technique is based on the Koopman framework, where the concept is to lift the nonlinear state space to a (possibly) infinite-dimensional space through the so-called observable functions. The dynamics of the original system are preserved and governed by a linear Koopman operator, enabling the representation of the system dynamics via a linear dynamical description [brunton_overview], [book_koopman] (for a more in depth-overview of the history of Koopman operator theory and the state-of-the-art see [Mezic_art_21]). In practice, by choosing a dictionary of a finite number of observables a priori to construct time-shifted data matrices, linear Koopman-based models have been commonly obtained using simple least squares estimation [Mauroy_sysid]. One such approach, called dynamic mode decomposition (DMD) [rowley_dmd], is based on constructing time-shifted data matrices using directly measured state variables associated with the system. If the dictionary consists of nonlinear functions of the state, this technique is known as extended DMD (EDMD) [williams_edmd]. However, among many challenges related to statistical consistency, availability of state-measurements, etc., the main difficulty with these powerful methods lies in choosing a finite number of lifting functions such that, in the lifted state-space, a linear time-invariant (LTI) model exists that can capture well the dynamic behavior of the original nonlinear system. While there exist methods for the automatic selection of the observables (see [Sindy], [yeung_nn]), they still rely on an a priori choice of a dictionary of functions, which many times are difficult to select and even characterize the resulting approximation error by them.
@@
To circumvent this, a viable approach has been found in learning the lifting functions from data by the use of machine learning methods such as Gaussian processes [Lian_gp], kernel-based methods [kernel_bounds], [kernel_edmd], or various forms of Artificial Neural Networks (ANNs) [lusch_nn, otto_bilinear, Otto, deepkoco_nn]. Due to their flexibility in describing multiple model structures, applicability to large datasets, many successful applications of these methods have been reported in the literature to obtain accurate and compact Koopman models in practice.
@@
However, a common drawback of learning-based methods together with the (E)DMD approaches is (i) the assumption of full-state measurement (e.g. [lusch_nn], [Otto]), which is rarely the case in engineering applications. Some works such as [Ketthong] and [yeung_nn] do address partial state observations, either by only lifting the output [yeung_nn] or by implementing a DMD version that uses time-delayed measurements [Ketthong]. In a different approach, [otto_bilinear] employs a Kalman filter to estimate the lifted state. Nevertheless, a systematic framework for addressing partial state measurements in the context of data-driven Koopman modelling is still lacking at large. Furthermore, despite the powerful capabilities of these approaches that have been demonstrated in multiple examples, generally (ii) little consideration of measurement or process noise is taken, which can lead to serious bias of the models when applied in real-world applications. Only a few papers present examples where measurement noise is even present in the data (e.g. [Jiang], [takeishi_noise]) and often only the robustness of the methods is analyzed (e.g. [Sinha]). While there are works that add process noise directly to the lifted representation (e.g. [otto_bilinear]), the way noise enters the Koopman model is merely an assumption. As such, ensuring statistical guarantees of consistency of the estimators remains an open question in the literature. A third important issue is that (iii) the estimation of Koopman models for systems with inputs has only recently been investigated, either through a nonlinear lifting [bonnert_nn] or by using state- and input-dependent observables, together with input increments [deepkoco_nn]. However, this often leads to models that have limited applicability, as it is more difficult to analyze dynamical aspects of the system or to design controllers to regulate the behavior compared to other model classes. Alternatively, due to their simple structure, multiple works assume a fully LTI Koopman model (e.g. [Korda], [Mamakoukas]) or, lately, bilinear (e.g. [Bruder, otto_bilinear, Schulze]). However, it is still largely unexplored how the approximation capability of these model structures in a finite dimensional setting compares to using more complex input structures such as control affine or nonlinear in both state and input, as given in [CDC_Id_Iacob, Aut_Iacob_inputs, Shi_koopman].
@@
To overcome challenges (i)-(iii), we introduce a flexible Koopman model learning method under control inputs, partial measurements, and with statistical guarantees of consistency under process and measurement noise. For this purpose, a deep-learning-based state-space encoder approach is proposed, which is implemented in the deepSI toolbox111deepSI toolbox available at https://github.com/MaartenSchoukens/deepSI in Python. The main advantages of the approach together with our contributions [^2] The present paper extends the conference paper [CDC_Id_Iacob] in terms of introducing an innovation noise structure in the Koopman model to handle process and measurement noise, proving the consistency of the estimator, studying various lifted structures for control inputs and providing extensive analysis and testing of the capabilities of the method on benchmarks and real-world data. are as follows:
@@ 2.1 Koopman embedding of nonlinear systems
with $x_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ being the state variable, $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{x}}}$ is a bounded nonlinear state transition map and $k\in\mathbb{Z}$ is the discrete time. The initial condition is denoted by $x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$ and we assume that $\mathbb{X}$ is forward invariant under $f(\cdot)$, i.e., $f(\mathbb{X})\subseteq\mathbb{X}$, see [Aut_Iacob_inputs]. The Koopman framework uses observable functions $\phi\in\mathcal{F}$ to lift the system (1) to a higher dimensional space with linear dynamics. These observables $\phi\mathrel{\mathop{\ordinarycolon}}\mathbb{X}\rightarrow\mathbb{R}$ are scalar functions (generally nonlinear) and are from a Banach function space $\mathcal{F}$. As described in [book_koopman], the Koopman operator $\mathcal{K}\mathrel{\mathop{\ordinarycolon}}\mathcal{F}\rightarrow\mathcal{F}$ associated with (1) is defined through:
@@
Although the Koopman framework typically requires $\mathcal{F}$ to be spanned by an infinite number of basis functions to fully describe the dynamics of (1), for practical use, an $n_{\mathrm{f}}$-dimensional linear subspace $\mathcal{F}_{n_{\mathrm{f}}}\subset\mathcal{F}$ is considered, with $\mathcal{F}_{n_{\mathrm{f}}}=\mathrm{span}\left\{\phi_{j}\right\}^{n_{\mathrm{f}}}_{j=1}$. The finite-dimensional approximation of the Koopman operator $\mathcal{K}$ can be described using the projection operator $\Pi\mathrel{\mathop{\ordinarycolon}}\mathcal{F}\rightarrow\mathcal{F}_{n_{\mathrm{f}}}$, and is given by:
(4) $\mathcal{K}_{n_{\mathrm{f}}}=\left.\Pi\mathcal{K}\right|_{\mathcal{F}_{n_{\mathrm{f}}}}\mathrel{\mathop{\ordinarycolon}}\mathcal{F}_{n_{\mathrm{f}}}\rightarrow\mathcal{F}_{n_{\mathrm{f}}}.$
In practice, the Koopman matrix representation $A\in\mathbb{R}^{n_{\mathrm{f}}\times n_{\mathrm{f}}}$ is commonly used [book_koopman]:
@@
Note that there exist classes of systems that admit an exact finite dimensional embedding and $\mathcal{K}_{n_{\mathrm{f}}}$ captures exactly the effect of $\mathcal{K}$. For example, [Iacob_poly_system] builds on the well known example discussed in [Brunton_2016], [Iacob_CT_block_embedding], or [otto_bilinear], and develops an exact embedding algorithm for network interconnections of block-oriented polynomial systems. Also, [Levine_86, Wang_23, Wong_83] discuss classes of nonlinear systems that admit an exact embedding via immersion. Next, we introduce the lifted state $z_{k}=\Phi(x_{k})$, where $\Phi(x_{k})=\begin{bmatrix}\phi_{1}(x_{k})&\cdots&\phi_{n_{\mathrm{f}}}(x_{k})\end{bmatrix}^{\top}$. The lifted finite dimensional linear representation of (1) is then given by:
@@
The main challenge of the Koopman framework is the selection of the observables, including their number, to obtain a suitable approximation in terms of an appropriate norm (or an exact embedding) of the nonlinear system [book_koopman]. Additionally, it is often not clearly stated in the literature that a linear system whose dynamics are governed by the Koopman matrix $A$ is only equivalent in terms of behavior (collections of all solution trajectories) to the original nonlinear system (1) if explicit nonlinear constraints are imposed on the initial condition of the lifted state, i.e., equivalent trajectories are only part of a manifold in the extended solution space. We explore this further through a simple example.
@@ 2.2 Linear representations subject to nonlinear constraints
To illustrate the concept, we consider the following polynomial system represented by (1), similar to the well-studied example described in [Brunton_2016]:
@@
where $a,b,c\in\mathbb{R}$ are constant parameters and $x_{k,i}$ denotes the $i^{\text{th}}$ element of $x_{k}$. By considering solutions of (7) only on $0,\infty)$ with initial condition $x_{0}\in\mathbb{R}^{2}$, the feasible trajectories are given by:
@@
(10) $\mathcal{B}_{\mathcal{K}}=\left\{z\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{+}_{0}\rightarrow\mathbb{R}^{3}\mid\text{s.t. }z_{k+1}=Az_{k}\right\}.$
Note that (10) represents an unrestricted LTI behavior. It is easy to show that $\Phi(\mathcal{B})\subseteq\mathcal{B}_{\mathcal{K}}$, as any $z_{k}\in\mathcal{B}_{\mathcal{K}}$ with $z_{0}\in\mathbb{R}^{3}$ for which $z_{0,3}\neq z^{2}_{0,1}$ will not correspond to a solution of (7), i.e., $\Phi^{-1}(z_{k})=x_{k}\notin\mathcal{B}$. By introducing the constraint $\Psi\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{3}\rightarrow\mathbb{R},\Psi(z_{k})=z^{2}_{k,1}-z_{k,3}$, the solution set (10) with constraint $\Psi$ is:
(11) $\hat{\mathcal{B}}_{\mathcal{K}}=\left\{z\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{+}_{0}\rightarrow\mathbb{R}^{3}\mid\text{s.t. }z_{k+1}=Az_{k},\;\Psi(z_{0})=0\right\}.$
Then, it is possible to show that $\Phi(\mathcal{B})=\hat{\mathcal{B}}_{\mathcal{K}}$ holds. To illustrate this, Fig. 1 shows the simulated trajectories of system (10) with $a=0.99$, $b=0.9$ and $c=0.9$ and the constraint $\Psi$, which we call the compliant surface. As can be seen in Fig. 1, only solutions (in green) starting on the compliant surface remain on the compliant surface and correspond to solutions (in black) of the original nonlinear system (7). This example highlights the need for additional constraints on the Koopman form, or, as we now call it, the embedding of (1), to guarantee a bijective relationship between the solution sets.
Note that, when $x_{0}$ is known and the observable set $\Phi$ is given, this nonlinear condition on the lifted states is alternatively defined by $z_{0}=\Phi(x_{0})$. This approach has been extensively used in the Koopman literature, in both theoretical and application oriented works, see e.g. [bonnert_nn, Cisneros20, Mauroy_sysid, surana_obs], and for a discussion on the connection between the nonlinear and lifted manifolds for systems of the type (7), one can consult [Brunton_2016]. However, in an identification setting where information on $x_{0}$ is not available or only partially available (in contrast to explicitly lifting the state via $\Phi(x_{0})$ as done in [Brunton_2016]), to construct a lifted model with the constrained solution set, one needs to include the $\Psi(z_{0})=0$ condition. Next, we explore observability and reconstructability of $z$ in view of this discussion.
@@ 2.3 Observability and reconstructability
where $h\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{y}}}$ is a nonlinear output map. Given $x_{0}\in\mathbb{R}^{n_{\mathrm{x}}}$, the output map for the state $x$ associated with the nonlinear dynamics represented by (1) and (12) is:
@@
where $h\circ_{2}f(x_{0})=h\circ f\circ f(x_{0})$ and $h\circ_{n}f(x_{0})=h\circ f\circ_{n-1}f(x_{0})$ for $n>2$. Let $y^{n-1}_{0}=\begin{bmatrix}y^{\top}_{0}&\cdots&y^{\top}_{n-1}\end{bmatrix}^{\top}$. As described in [nl_obs_1982], the representation satisfies the so-called observability rank condition at $x_{0}$ if, for $n=n_{\mathrm{x}}$, the rank of the Jacobian of $\mathcal{O}_{\mathrm{x},n}$ at $x_{0}$ is $n_{\mathrm{x}}$. If this condition is met, the representation is strongly locally observable at $\mathbb{X}_{0}$, where $\mathbb{X}_{0}$ is a neighborhood of $x_{0}$, and $\mathcal{O}_{\mathrm{x},n}$ is a diffeomorphism, i.e., it is invertible, on $\mathbb{X}_{0}$ [Isidori]. We denote its inverse as $\Lambda_{\mathrm{x},n}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{nn_{\mathrm{y}}}\rightarrow\mathbb{X}_{0}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$, such that
@@
Throughout the paper, we call $\Lambda_{\mathrm{x},n}$ the observability map, which can be used to compute the initial condition $x_{0}$ from future values of the output $y$. Conversely, the reconstructability concept is used to calculate the initial condition from past values of the output. For $n\geq 1$, it holds that
@@
(16) $x_{0}=\circ_{n-1}f(\Lambda_{\mathrm{x},n}(y^{0}_{-n+1}))=\mathrel{\mathop{\ordinarycolon}}\Pi_{\mathrm{x},n}(y^{0}_{-n+1})$
for all $x_{0}\in\mathbb{X}_{0}$. The function $\Pi_{\mathrm{x},n}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{nn_{\mathrm{y}}}\rightarrow\mathbb{X}_{0}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$ is called the reconstructability map.
@@
(17) $\begin{bmatrix}y_{0}\\ y_{1}\\ \vdots\\ y_{n-1}\\ 0\end{bmatrix}=\begin{bmatrix}\begin{pmatrix}C\\ CA\\ \vdots\\ CA^{n-1}\end{pmatrix}z_{0}\\ \Psi(z_{0})\end{bmatrix}\mathrel{\mathop{\ordinarycolon}}=\begin{bmatrix}\mathcal{O}^{\text{lin}}_{\mathrm{z},n}z_{0}\\ \Psi(z_{0})\end{bmatrix}\mathrel{\mathop{\ordinarycolon}}=\mathcal{O}_{\mathrm{z},n}(z_{0}).$
In an LTI sense, the lifted system representation would be observable if there is an $n\geq 1$ such that $\text{rank}(\mathcal{O}^{\text{lin}}_{\mathrm{z},n})=n_{\mathrm{z}}$. However, as observed in Section 2.2, it is also necessary to consider the nonlinear constraints $\Psi\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n_{\mathrm{z}}}\rightarrow\mathbb{R}^{n_{\mathrm{c}}}$ to ensure compliance of the initial condition $z_{0}$. Hence, if there is an $n\geq 1$ such that the Jacobian of $\mathcal{O}_{\mathrm{z},n}(z_{0})$ has full rank $n_{\mathrm{z}}$, which implies that the mapping $\mathcal{O}_{\mathrm{z},n}$ is locally invertible on a neighborhood $\mathbb{Z}_{0}$ of $z_{0}$, then $z_{0}$ is uniquely determined from $y^{n-1}_{0}$ and the constraint $\Psi(z_{0})$. Then, there exists a $\Lambda_{\mathrm{z},n}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{nn_{\mathrm{y}}}\rightarrow\mathbb{Z}_{0}\subseteq\mathbb{R}^{n_{\mathrm{z}}}$, such that